Assuming constant pressure, rank these reactions from most energy released by the system to most energy absorbed by the system, based on the following descriptions:

Surroundings get colder and the system decreases in volume.

Surroundings get hotter and the system expands in volume.

Surroundings get hotter and the system decreases in volume.

Surroundings get hotter and the system does not change in volume.

Also assume that the magnitude of the volume and temperature changes are similar among the reactions.

Rank from most energy released to most energy absorbed. To rank items as equivalent, overlap them.

Part B

A mole of X reacts at a constant pressure of 43.0 atm via the reaction

X(g)+4Y(g)→2Z(g),      ΔH∘=−75.0 kJ

Before the reaction, the volume of the gaseous mixture was 5.00 L. After the reaction, the volume was 2.00 L. Calculate the value of the total energy change, ΔE, in kilojoules.

Express your answer with the appropriate units.

ΔE =
Answer #1
Concepts and reason

System and surroundings are the two terms in thermodynamics that are the parts of the universe. Heat transfer takes place in between the system and its surroundings.

Total energy of the system at constant pressure depends upon the heat absorbed by the system and expansion work done by the system.

Fundamentals

If the surroundings get cooler, it means the system absorbs the energy from surroundings.

If the surroundings get hotter, it means the system releases the energy to the surroundings.

Volume decrease means the system absorbs energy; that means work is done on the system.

Volume increase means the system releases energy and work is done by the system.

According to the first law of thermodynamics, the energy of the system is defined as the sum of the heat involved in the system and work done the system.

ΔE=q+WΔE=ΔHPΔV\begin{array}{l}\\{\rm{\Delta E}}\,{\rm{ = }}\,{\rm{q + W}}\\\\{\rm{\Delta E}}\,{\rm{ = }}\,{\rm{\Delta H - P\Delta V}}\\\end{array}

Surroundings get hotter and the system expands in volume is the most energy-releasing process.

Surroundings get hotter and the system does not change in volume is the second most energy-releasing process.

Surroundings get hotter and the system decrease in volume is the third most energy-releasing process in the given set.

Surroundings get colder and the system decreases in volume is the least energy-releasing process in the given set.

Therefore, the order is

1.Surroundings get hotter and the system expands in volume

2.Surroundings get hotter and the system does not change in volume

3.Surroundings get hotter and the system decreases in volume

4.Surroundings get colder and the system decreases in volume

(B)

Fromthefirstlawofthermodynamics,ΔE=q+WΔE=ΔHPΔVGiven,ΔH=75.0kJP=43.0atmΔV=2L5L=3LΔE=75kJ(43.0atm×3L×101.325JLatm)=61.9kJ\begin{array}{c}\\{\rm{From}}\,{\rm{the}}\,{\rm{first}}\,{\rm{law}}\,{\rm{of}}\,{\rm{thermodynamics,}}\\\\{\rm{\Delta E}}\,{\rm{ = }}\,{\rm{q}}\,{\rm{ + }}\,{\rm{W}}\\\\{\rm{\Delta E}}\,{\rm{ = }}\,{\rm{\Delta H - P\Delta V}}\\\\{\rm{Given,}}\\\\{\rm{\Delta H}}\,{\rm{ = }}\,{\rm{ - 75}}{\rm{.0}}\,{\rm{kJ}}\\\\{\rm{P = }}\,{\rm{43}}{\rm{.0atm}}\\\\{\rm{\Delta V}}\,{\rm{ = }}\,{\rm{2L - 5L}}\,{\rm{ = - 3L}}\\\\{\rm{\Delta E}}\,{\rm{ = }}\,{\rm{ - 75kJ - }}\,\left( {{\rm{43}}{\rm{.0atm \times - 3L \times }}\frac{{{\rm{101}}{\rm{.325J}}}}{{{\rm{L}}\,{\rm{atm}}}}} \right)\\\\\,\,\,\,{\rm{ = }}\,{\rm{ - 61}}{\rm{.9kJ}}\\\end{array}

Ans:

Surroundings get hotter and the system expands in volume>Surroundings get hotter and the system does not change in volume >Surroundings get hotter and the system decreases in volume >Surroundings get colder and the system decreases in volume